# latexOverwriting file

### v1

@@ 2.2 Well-posedness for CBO and CBS
In this section, we state well-posedness results for CBO and CBS, referring for the proofs to Section 4. While for CBO, the following theorems are only generalizations of already existing results [10], the well-posedness proofs for the CBS particle and mean-field systems are new.
@@
In the case of CBO, Theorem 2.2 was already proved in [10, Theorem 2.1]. We now present well-posedness results for the mean-field SDEs, beginning with CBO. For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [10, Theorems 3.1 and 3.2].
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant-1$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.3) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
[…]
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant-1$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.5) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
@@ 2.3 Key preparatory lemma
We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of a standard non-asymptotic version of the central limit theorem known as Bikelis’ theorem, see for example [28, Theorem 3.7].
@@
Let $(Z_{j})_{j\in\mathbf{N}^{+}}$ be a family of i.i.d. $\mathbf{R}$-valued random variables such that $\mathbf{E}\left[\left\lvert Z_{1}\right\rvert^{3}\right]<\infty$. Then for all $R>\mathbf{E}\bigl{[}\left\lvert Z_{1}\right\rvert\bigr{]}$, there exists a constant $C>0$ such that
$\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant\frac{C}{J^{2}}.$
@@
Denote by $m$ and $\sigma^{2}$ the mean and variance of $Z_{j}$, respectively. Bikelis’ theorem provides the following non-asymptotic error estimate:
$\forall z\in\mathbf{R},\qquad\bigl{\lvert}\phi_{J}(z)-\phi(z)\bigr{\rvert}\leqslant\frac{1}{\sqrt{J}}\,\left(\frac{\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]}{\sigma^{3}(1+\lvert z\rvert)^{3}}\right),\qquad\phi_{J}(z):=\mathbf{P}\left[\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\frac{Z_{j}-m}{\sigma}\leqslant z\right],$
where $\phi$ is the cumulative distribution function of the $\mathcal{N}(0,1)$ Gaussian law. Thus, if $R>m$, we obtain
$\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]$ $=\mathbf{P}\left[\frac{1}{\sqrt{J}}\sum_{j=1}^{J}\frac{Z_{j}-m}{\sigma}\geqslant\sqrt{J}\left(\frac{R-m}{\sigma}\right)\right]$
$=1-\phi_{J}\left(\sqrt{J}k\right),\qquad k:=\frac{R-m}{\sigma}$
$\leqslant 1-\phi\left(\sqrt{J}k\right)+\frac{1}{\sqrt{J}}\,\left(\frac{\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]}{\sigma^{3}(1+\sqrt{J}k)^{3}}\right)=\mathcal{O}(J^{-2}),$
which completes the proof. ∎
@@ 2.4 Quantitative mean-field limit for CBO
We now present and prove the mean-field results, starting with CBO. Synchronous coupling for McKean-Vlasov diffusions with drift $b\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d}$ and diffusion $\sigma\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d\times d}$ assumes the following setting [14, 15]. Given i.i.d. Brownian motions $\bigl{(}W^{(j)}\bigr{)}_{j\in\mathbf{N}^{+}}$ and initial conditions $\bigl{(}X^{(j)}_{0}\bigr{)}_{j\in\mathbf{N}^{+}}$ sampled i.i.d. from some $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$, we consider for each $J\in\mathbf{N}^{+}$ the interacting particle system
$\forall j\in,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
@@
$\forall N\in\mathbf{N}^{+},\forall j\in,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$, and consider the systems (2.8) and (2.9) with the coefficients given by (2.10). Then for all $p\geqslant 2$ such that
$\frac{q}{3}\geqslant p\geqslant p_{\mathcal{M}}(s,\ell),$ (2.11)
where $p_{\mathcal{M}}$ is given in (2.3), there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\min\left\{\frac{p}{2},2-\frac{2p}{q}\right\}}.$
@@
In particular, we obtain a decay bound of order $\mathcal{O}\left(\frac{1}{J}\right)$ if $p=2$. If $p\geqslant 4$, then for each $\varepsilon>0$, decay of order $\mathcal{O}\left(\frac{1}{J^{2-\varepsilon}}\right)$ holds provided that $\overline{\rho}_{0}\in\mathcal{P}_{q}(^{d})$ with $q\geqslant\frac{2p}{\varepsilon}\vee 3p$.
@@
$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{2}{p}-\frac{2}{q}\right\}}.$
@@
See [14, Definitions 3.1 and 3.5] for more details.
@@
Fix $j\in$. As in Sznitman’s argument [15, Theorem 3.1], we begin by writing
@@
By the Burkholder–Davis–Gundy inequality, see [44, Theorem 7.3], it holds for any $t\in[0,T]$ that
@@
In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [15, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
@@
$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}$
Since $q\geqslant 3p$, Theorem 2.3 implies that $\mathbf{E}\left[\lvert Z_{j}\rvert^{3}\right]<\infty.$ Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
$\forall t\in[0,T],\qquad\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\Omega_{J,t}\right]\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant\frac{C}{J^{2}}.$
We can split the second term on the right-hand of (2.14) as follows:
@@
Let us now bound each term on the right-hand side of 2.16 separately.
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{2(q-p)}{q}}.$
@@
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{2(q-p)}{q}},$
@@
$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{2(q-p)}{q}}.$

### v2

@@ 2.2 Well-posedness for CBO and CBS
In this section, we state well-posedness results for CBO and CBS, referring for the proofs to Section 4. While for CBO, the following theorems are only generalizations of already existing results [carrillo2018analytical], the well-posedness proofs for the CBS particle and mean-field systems are new.
@@
In the case of CBO, Theorem 2.2 was already proved in [carrillo2018analytical, Theorem 2.1]. We now present well-posedness results for the mean-field SDEs, beginning with CBO. For $f\in\mathcal{A}(1,0,0)\cup A(1,2,2)$, the following result generalizes [carrillo2018analytical, Theorems 3.1 and 3.2].
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{M}}(s,\ell)$, Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.3) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
$\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert\overline{X}_{t}\right\rvert^{p}\right]<\infty\qquad\text{and}\qquad\sup_{t\in[0,T]}\Bigl{\lvert}\mathcal{M}_{\beta}(\overline{\rho}_{t})\Bigr{\rvert}<\infty,\qquad\overline{\rho}_{t}:=\operatorname{Law}(\overline{X}_{t}),$ (2.4)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with parameters $s\geqslant 0$, $\ell\geqslant 0$. Assume also that $\overline{\rho}_{0}\in\mathcal{P}_{p}(\mathbf{R}^{d})$ for $p\geqslant 2\vee p_{\mathcal{C}}(s,\ell)$, that $C_{\beta}(\overline{\rho}_{0})\succ 0$, and fix $x_{0}\sim\overline{\rho}_{0}$. Then, for each final time $T>0$, there exists a unique process $\overline{X}\in C([0,T],\mathbf{R}^{d})$ satisfying (1.5) in the strong sense with initial condition $\overline{X}_{0}=x_{0}$. Furthermore, it holds that
@@ 2.3 Key preparatory lemma
We first show a key preparatory lemma which is needed for the mean-field results. This result is a simple corollary of Markov’s inequality,
@@
Let $(Z_{j})_{j\in\mathbf{N}^{+}}$ be a family of i.i.d. -valued random variables such that $\mathbf{E}\left[\left\lvert Z_{1}\right\rvert^{r}\right]<\infty$ for some $r\geqslant 2$. Then for all $R>\mathbf{E}\bigl{[}\left\lvert Z_{1}\right\rvert\bigr{]}$, there exists a constant $C>0$ such that
$\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant CJ^{-\frac{r}{2}}.$
@@
This follows from a generalization of Chebychev’s inequality [MR3674428, Exercise 3.21]. Let
$X=\frac{1}{J}\sum_{j=1}^{J}Z_{j}.$
By the classical Marcinkiewicz–Zygmund inequality, it holds that
$\mathbf{E}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}$ $\leqslant J^{-r}C_{\rm MZ}(r)\mathbf{E}\left[\left(\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{2}\right)^{\frac{r}{2}}\right]$
$\leqslant J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\frac{1}{J}\sum_{j=1}^{J}\Bigl{\lvert}Z_{i}-\mathbf{E}[Z_{i}]\Bigr{\rvert}^{r}\right]=J^{-\frac{r}{2}}C_{\rm MZ}(r)\mathbf{E}\left[\Bigl{\lvert}Z_{1}-\mathbf{E}[Z_{1}]\Bigr{\rvert}^{r}\right],$
where we used the convexity of the function $x\mapsto x^{\frac{r}{2}}$ in the second inequality. Therefore, using the Markov inequality, we deduce that
$\mathbf{P}\left[X\geqslant R\right]\leqslant\mathbf{P}\Bigl{[}\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}\geqslant\bigl{(}R-\mathbf{E}[X]\bigr{)}^{r}\Bigr{]}$ $\leqslant\mathbf{E}\left[\frac{\bigl{\lvert}X-\mathbf{E}[X]\bigr{\rvert}^{r}}{\bigl{(}R-\mathbf{E}[X]\bigr{)}^{r}}\right]\leqslant\frac{CJ^{-\frac{r}{2}}}{\bigl{(}R-\mathbf{E}[X]\bigr{)}^{r}},$
which concludes the proof. ∎
@@ 2.4 Quantitative mean-field limit for CBO
We now present and prove the mean-field results, starting with CBO. Synchronous coupling for McKean-Vlasov diffusions with drift $b\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d}$ and diffusion $\sigma\colon\mathbf{R}^{d}\times\mathcal{P}(\mathbf{R}^{d})\rightarrow\mathbf{R}^{d\times d}$ assumes the following setting [ReviewChaintronI, ReviewChaintronII]. Given i.i.d. Brownian motions $\bigl{(}W^{(j)}\bigr{)}_{j\in\mathbf{N}^{+}}$ and initial conditions $\bigl{(}X^{(j)}_{0}\bigr{)}_{j\in\mathbf{N}^{+}}$ sampled i.i.d. from some $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$, we consider for each $J\in\mathbf{N}^{+}$ the interacting particle system
$\forall j\in\llbracket 1,J\rrbracket,\qquad X^{(j)}_{t}=X^{(j)}_{0}+\int_{0}^{t}b\left(X^{(j)}_{s},\mu^{J}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(X^{(j)}_{s},\mu^{J}_{s}\right)\mathrm{d}W^{(j)}_{s},$ (2.8)
@@
$\forall N\in\mathbf{N}^{+},\forall j\in\llbracket 1,J\rrbracket,\qquad\left\{\begin{aligned} \overline{X}^{(j)}_{t}&=X^{(j)}_{0}+\int_{0}^{t}b\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\left(\overline{X}^{(j)}_{s},\overline{\rho}_{s}\right)\,\mathrm{d}W^{(j)}_{s},\\ \overline{\rho}_{t}&=\operatorname{Law}\left(\overline{X}^{(j)}_{t}\right),\end{aligned}\right.$ (2.9)
@@
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ has moments of sufficiently high order. Consider the systems (2.8) and (2.9) with the coefficients given by (2.10). Then, for each $p>0$, there exists a constant $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\frac{1}{2}}.$ (2.11)
@@
We prove below that the following sharper statement holds. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$, and consider the systems (2.8) and (2.9) with the coefficients given by (2.10). If $q\geqslant\max\bigl{\{}4,2p_{\mathcal{M}}(s,\ell)\bigr{\}}$, where $p_{\mathcal{M}}$ is given in (2.3), then for all $p\in(0,\frac{q}{2}]$ there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-p}{2p^{2}},\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}.$ (2.12)
@@
$W_{p}\left(\rho^{J}_{[0,T]},\overline{\rho}^{\otimes J}_{[0,T]}\right)\leqslant CJ^{-\frac{1}{2}}.$
@@
See [ReviewChaintronI, Definitions 3.1 and 3.5] for more details.
@@
As mentioned in Remark 2.7, we prove the sharp result (2.12). It is sufficient to prove this statement for all $p\in[2\vee p_{\mathcal{M}},\frac{q}{2}]$ (by the assumption on $q$, this interval is nonempty). Indeed, if the statement holds true in this case, then by Jensen’s inequality, it holds for all $\mathfrak{p}\in(0,2\vee p_{\mathcal{M}}]$ that
$\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{\mathfrak{p}}\right]\right)^{\frac{1}{\mathfrak{p}}}$ $\leqslant\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{2\vee p_{\mathcal{M}}}\right]\right)^{\frac{1}{2\vee p_{\mathcal{M}}}}$
$\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}}=CJ^{-\min\left\{\frac{1}{2},\frac{q-\mathfrak{p}}{2\mathfrak{p}^{2}},\frac{q-2\vee p_{\mathcal{M}}}{2(2\vee p_{\mathcal{M}})^{2}}\right\}},$
where we used that the function $z\mapsto(q-z)/z^{2}$ is decreasing. Take $p\in[2\vee p_{\mathcal{M}},q/2]$ and fix $j\in\llbracket 1,J\rrbracket$. As in Sznitman’s argument [ReviewChaintronII, Theorem 3.1], we begin by writing
@@
By the Burkholder–Davis–Gundy inequality, see [MR2380366, Theorem 7.3], it holds for any $t\in[0,T]$ that
@@
In order to deal with the fact that the weighted mean does not have global Lipschitz properties in the sense of McKean’s framework [ReviewChaintronII, Theorem 3.1], we introduce for each $t\in[0,T]$ and $J\in\mathbf{N}^{+}$ the set
@@
$Z_{j}=\sup_{t\in[0,T]}\left\lvert\overline{X}^{(j)}_{t}\right\rvert^{p}.$
By Theorem 2.3 it holds that $\mathbf{E}\left[\lvert Z_{j}\rvert^{q/p}\right]<\infty.$ Note that $q\geqslant 2p$. Therefore, by Lemma 2.5 there exists a constant $C>0$ such that
$\forall t\in[0,T],\qquad\forall J\in\mathbf{N}^{+},\qquad\mathbf{P}\left[\Omega_{J,t}\right]\leqslant\mathbf{P}\left[\frac{1}{J}\sum_{j=1}^{J}Z_{j}\geqslant R\right]\leqslant CJ^{-\frac{q}{2p}}.$
We split the second term on the right-hand of (2.15) as follows:
@@
Let us now bound each term on the right-hand side of (2.17) separately.
@@
    $\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\mathsf{1}_{\Omega_{J,s}}\right]\leqslant\left(\mathbf{E}\left[\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{q}\right]\right)^{\frac{p}{q}}\left(\mathbf{P}\left[\Omega_{J,s}\right]\right)^{{}^{\frac{q-p}{q}}}\leqslant 2^{p}\kappa^{\frac{p}{q}}J^{-\frac{q-p}{2p}}.$
@@
$\mathbf{E}\left\lvert\mathcal{M}_{\beta}\left(\mu^{J}_{s}\right)-\mathcal{M}_{\beta}\left(\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}},$
@@
$\mathbf{E}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\mu}^{J}_{s}\right)\right\rvert^{p}\leqslant C\mathbf{E}\left\lvert X^{(j)}_{s}-\overline{X}^{(j)}_{s}\right\rvert^{p}+CJ^{-\frac{q-p}{2p}}.$